# Characterizing Graph Datasets for Node Classification: Homophily–Heterophily Dichotomy and Beyond

### v4

@@ 1 Introduction
Graphs are a natural way to represent data from various domains such as social networks, citation networks, molecules, transportation networks, text, code, and others. Machine learning on graph-structured data has experienced significant growth in recent years, with Graph Neural Networks (GNNs) showing particularly strong results. Many variants of GNNs have been proposed Kipf and Welling (2017); Hamilton et al. (2017); Veličković et al. (2018); Xu et al. (2019), most of them can be unified by a general Message Passing Neural Networks (MPNNs) framework Gilmer et al. (2017). MPNNs combine node features (attributes) with graph topology to learn complex dependencies between the nodes. For this, MPNNs iteratively update the representation of each node by aggregating information from the previous-layer representations of the node itself and its neighbors.
In many real-world networks, edges tend to connect similar nodes: users in social networks tend to connect to users with similar interests, and papers in citation networks mostly cite works from the same research area. This property is usually called homophily. The opposite of homophily is heterophily: for instance, in social networks, fraudsters rarely connect to other fraudsters, while in dating networks, edges often connect the opposite genders. Early research on GNNs mainly focused on homophilous graphs, and recently there have been discussions whether specialized models are needed for the heterophilous setting Abu-El-Haija et al. (2019); Pei et al. (2020); Zhu et al. (2020); Ma et al. (2022); Luan et al. (2022); Platonov et al. (2023).
To measure the level of homophily, several homophily measures are used in the literature Abu-El-Haija et al. (2019); Lim et al. (2021); Pei et al. (2020); Zhu et al. (2020), but these measures may significantly disagree with each other. In this work, we start by addressing the problem of how to properly measure the homophily level of a graph. For this, we formalize some desirable properties of a reasonable homophily measure and check which measures satisfy which properties. One essential property is called constant baseline and, informally speaking, it requires a measure to be not biased towards particular numbers of classes or their size balance. Our analysis reveals that commonly used homophily measures do not satisfy this property and thus cannot be compared across different datasets. In contrast, a measure that we call adjusted homophily (a.k.a. assortativity coefficient) satisfies most of the desirable properties while being rarely used in graph ML literature. Based on our theoretical analysis, we advise using adjusted homophily as a better alternative to the commonly used measures.
@@ 2.1 Popular homophily measures
Many GNN models implicitly make a so-called homophily assumption: that similar nodes are connected. Similarity can be considered in terms of node features or node labels. Usually, label homophily is analyzed, and we also focus on this aspect, leaving feature homophily for further studies. There are several commonly used homophily measures in the literature. Edge homophily Abu-El-Haija et al. (2019); Zhu et al. (2020) is the fraction of edges that connect nodes of the same class:
@@
Node homophily Pei et al. (2020) computes the fraction of neighbors that have the same class for all nodes and then averages these values across the nodes:
@@
These two measures are intuitive but have the downside of being sensitive to the number of classes and their balance, which makes them hard to interpret and incomparable across different datasets Lim et al. (2021). For example, suppose that each node in a graph is connected to one node of each class. Then, both edge homophily and node homophily for this graph will be equal to $\frac{1}{C}$. Thus, these metrics will produce widely different values for graphs with different numbers of classes, despite these graphs being similar in exhibiting no homophily. To fix these issues, Lim et al. (2021) propose another homophily measure sometimes referred to as class homophily Luan et al. (2022). Class homophily measures excess homophily compared to a null model where edges are independent of the labels. More formally,
@@ 2.2 Desirable properties for homophily measures
Above, we discussed some disadvantages of existing homophily measures. In this section, we formalize and extend this discussion: we propose a list of properties desirable for a good homophily measure. Our analysis is motivated by recent studies of clustering and classification performance measures Gösgens et al. (2021a, b), but not all their properties can be transferred to homophily measures. For instance, we do not require symmetry — a property that a measure does not change when we swap the compared objects — since homophily compares entities of different nature (a graph and a labeling). For the same reason, the distance property (requiring a measure to be linearly transformed to a metric distance) cannot be defined. On the other hand, some of our properties are novel.
@@
For instance, edge homophily and node homophily are empty class tolerant, while class homophily is not. Empty class tolerance is a new property that was not discussed in Gösgens et al. (2021a, b) since classification and clustering evaluation measures are used within one given dataset, see Appendix B.3 for more details.
@@
In contrast to Gösgens et al. (2021a, b), our notion of monotonicity requires the property to hold across graphs with different numbers of classes. This is caused by the empty class tolerance property.
@@ 2.4 Adjusted homophily
where we use the notation $\bar{p}(k)=\frac{D_{k}}{2|E|}$ (degree-weighted distribution of class labels). This measure is known in graph analysis literature as assortativity coefficient Newman (2003). While assortativity is a general concept that is often applied to node degrees, it reduces to (2) when applied to categorical node attributes on undirected graphs. Unfortunately, this measure is rarely used in graph ML literature (Suresh et al. (2021) is the only work we are aware of that uses it for measuring homophily), while our theoretical analysis shows that it satisfies many desirable properties. Indeed, the following theorem holds (see Appendix B.4 for the proof).
@@ 2.5 Edge-wise homophily vs classification evaluation measures
We conclude the analysis of homophily measures by establishing a connection between them and classification evaluation measures Gösgens et al. (2021a). For this, let us first define edge-wise homophily measures. We say that a homophily measure is edge-wise if it is a function of the class adjacency matrix that we now define. Since we consider undirected graphs, each edge $\{u,v\}\in E$ gives two ordered pairs of nodes $(u,v)$ and $(v,u)$. We can define a class adjacency matrix $\mathcal{C}$ as follows: each matrix element $c_{ij}$ denotes the number of edges $(u,v)$ such that $y_{u}=i$ and $y_{v}=j$. Since the graph is undirected, the matrix $\mathcal{C}$ is symmetric. Note that monotonicity can be naturally put in terms of the class adjacency matrix: adding an edge between two nodes of the same class $i$ corresponds to incrementing the diagonal element $c_{ii}$ by two, and adding an edge between two nodes of different classes $i$ and $j$ corresponds to incrementing $c_{ij}$ and $c_{ji}$ by one.
@@
Interestingly, both Cohen’s Kappa and Matthews correlation coefficient correspond to adjusted homophily. As argued in Gösgens et al. (2021a), the Matthews coefficient is one of the best classification evaluation measures in terms of its theoretical properties. Our extended analysis confirms this conclusion for the corresponding homophily measure: we prove a stronger version of asymptotic constant baseline and also establish a stronger variant of monotonicity in the interval of large values of $h_{adj}$. The latter result is essential since in Gösgens et al. (2021a) it was only claimed that monotonicity is violated when $C>2$.
Another measure advised by Gösgens et al. (2021a) is symmetric balanced accuracy. Since in our case the class adjacency matrix is symmetric, it gives the same expression as balanced accuracy: $h_{bal}=\frac{1}{C}\sum_{k=1}^{C}\frac{|(u,v):y_{u}=y_{v}=k|}{D_{k}}.$ The obtained measure satisfies the maximal and minimal agreement properties. However, it is not empty class tolerant and thus is not monotone. The asymptotic constant baseline is also not satisfied: the value $c_{base}=1/C$ depends on the number of classes. Thus, despite this measure being suitable for classification evaluation, it cannot be used as a homophily measure. This difference is caused by the fact that homophily measures have to be comparable across datasets with different numbers of classes, while for classification evaluation it is not required.
@@
Finally, we also note that homophily measures can be directly related to community detection evaluation measures including the well-established characteristic in graph community detection literature called modularity Newman and Girvan (2004). See Appendix B.5 for a detailed discussion.
@@ 3 Label informativeness
To give an example, among strictly heterophilous graphs in which nodes never connect to nodes of the same class, there can be those where edges are drawn between particular pairs of classes (Figure 1) and those where the class of a node cannot be derived from the class of its neighbor (Figure 1). While adjusted homophily correctly captures the absence of homophily in these graphs, it is not designed to identify which type they belong to. However, distinguishing such graphs is practically important: informative neighbors can be very useful for models accounting for the graph structure.
We define a characteristic measuring the informativeness of a neighbor’s label for a node’s label. For example, in Figure 1, the neighbor’s label uniquely defines the node’s label. Thus, the node classification task is simple on this dataset, and we want our informativeness to be maximal for such graphs. Let us formalize this idea. Assume that we sample an edge $(\xi,\eta)\in E$ (from some distribution). The class labels of nodes $\xi$ and $\eta$ are then random variables $y_{\xi}$ and $y_{\eta}$. We want to measure the amount of knowledge the label $y_{\eta}$ gives for predicting $y_{\xi}$. The entropy $H(y_{\xi})$ measures the ‘hardness’ of predicting the label of $\xi$ without knowing $y_{\eta}$. Given $y_{\eta}$, this value is reduced to the conditional entropy $H(y_{\xi}|y_{\eta})$. In other words, $y_{\eta}$ reveals $I(y_{\xi},y_{\eta})=H(y_{\xi})-H(y_{\xi}|y_{\eta})$ information about the label. To make the obtained quantity comparable across different datasets, we say that label informativeness is the normalized mutual information of $y_{\xi}$ and $y_{\eta}$:
@@ 4.1 Characterizing real graph datasets
It is expected that datasets with high homophily $h_{adj}$ also have high $\mathrm{LI}$ since homophily implies informative neighbor classes. For medium-level homophily, $\mathrm{LI}$ can behave differently: for instance, ogbn-arxiv and twitter-hate have similar homophily levels, while the neighbors in ogbn-arxiv are significantly more informative. For heterophilous datasets, $\mathrm{LI}$ can potentially be very different, as we demonstrate in Section 4.2 on synthetic and semi-synthetic data. However, most existing heterophilous datasets have $\mathrm{LI}\approx 0$. This issue is partially addressed by new heterophilous datasets recently proposed in Platonov et al. (2023). For the proposed roman-empire dataset, $\mathrm{LI}=0.11$ and $h_{adj}=-0.05$. Thus, while being heterophilous, this dataset has non-zero label informativeness, meaning that neighboring classes are somewhat informative. We believe that datasets with more interesting connectivity patterns will be collected in the future.
@@ 4.2 Correlation of LI with GNN performance
Recently, it has been shown that standard GNNs can sometimes perform well on non-homophilous datasets Luan et al. (2022); Ma et al. (2022). We hypothesize that GNNs can learn more complex relationships between nodes than just homophily, and they will perform well as long as the node’s neighbors provide some information about this node. Thus, we expect $\mathrm{LI}$ to better correlate with the performance of GNNs than homophily. To illustrate this, we use carefully designed synthetic and semi-synthetic data. First, it allows us to cover all combinations of homophily levels and label informativeness. Second, we can control that only a connection pattern changes while other factors affecting the performance are fixed.
@@
To start with the most simple and controllable setting, we generate synthetic graphs via a variant of the stochastic block model (SBM) Holland et al. (1983). In this model, the nodes are divided into $C$ clusters, and for each pair of nodes $i,j$, we draw an edge between them with probability $p_{c(i),c(j)}$ independently of all other edges. Here $c(i)$ is a cluster assignment for a node $i$, which in our case corresponds to the node label $y_{i}$.
@@
We generated graphs according to the procedure described above with the expected node degree of 10 and various combinations of $p_{0},p_{1},p_{2}$. Given the class labels, the features are taken from the four largest classes in the cora dataset Sen et al. (2008); Namata et al. (2012); Yang et al. (2016); McCallum et al. (2000). We use the obtained graphs to train four popular GNN models: GCN Kipf and Welling (2017), GraphSAGE Hamilton et al. (2017), GAT Veličković et al. (2018), and Graph Transformer (GT) Shi et al. (2020). In total, we run more than 20000 experiments on synthetic datasets for each of the four models. A detailed description of data generation, training setup, and hyperparameters used is provided in Appendix F.1.
@@
Figure 2 shows the results for GraphSAGE. It can be seen that the performance is much more correlated with $\mathrm{LI}$ than with homophily. In particular, when $\mathrm{LI}$ is high, GraphSAGE achieves good performance even on strongly heterophilous graphs with negative homophily. We refer to Appendix F.2 for additional visualizations.
[…]
@@
Semi-synthetic data from Ma et al. (2022)
Ma et al. (2022) also argue that standard GNNs can perform well on certain heterophilous graphs. They construct semi-synthetic graphs by adding inter-class edges following different patterns to several real-world graphs, thus obtaining several sets of graphs with varying levels of homophily. Ma et al. (2022) run experiments on these graphs and note that a standard GNN achieves strong performance on some heterophilous graphs.
Table 4: Spearman correlation between model accuracy and characteristics of semi-synthetic datasets from Ma et al. (2022)
@@
We train GCN, GraphSAGE, GAT, and GT on the same modifications of the cora and citeseer graphs used in Ma et al. (2022) and find that the models achieve strong performance when the graphs have high label informativeness. Training setup and hyperparameters used in these experiments are the same as above and are described in Appendix F.1. The performance of GraphSAGE on cora is shown in Figure 2. In Appendix F.3, we also show the performance of GraphSAGE on the citeseer dataset. Table 4 shows the Spearman correlation coefficients between accuracy and various homophily measures or $\mathrm{LI}$ for all models and both datasets. The correlation coefficients for homophily measures are all negative, while for $\mathrm{LI}$ they are positive and sufficiently large. This again confirms that $\mathrm{LI}$ indicates whether graph structure is helpful for GNNs.
Additionally, we conduct experiments with other datasets: Appendix F.4 describes the results for the LFR benchmark Lancichinetti et al. (2008) and Appendix F.5 presents an experiment on synthetic data from Luan et al. (2022).

### v5

@@ 1 Introduction
Graphs are a natural way to represent data from various domains such as social networks, citation networks, molecules, transportation networks, text, code, and others. Machine learning on graph-structured data has experienced significant growth in recent years, with Graph Neural Networks (GNNs) showing particularly strong results. Many variants of GNNs have been proposed [16, 10, 43, 45], most of them can be unified by a general Message Passing Neural Networks (MPNNs) framework [7]. MPNNs combine node features (attributes) with graph topology to learn complex dependencies between the nodes. For this, MPNNs iteratively update the representation of each node by aggregating information from the previous-layer representations of the node itself and its neighbors.
In many real-world networks, edges tend to connect similar nodes: users in social networks tend to connect to users with similar interests, and papers in citation networks mostly cite works from the same research area. This property is usually called homophily. The opposite of homophily is heterophily: for instance, in social networks, fraudsters rarely connect to other fraudsters, while in dating networks, edges often connect the opposite genders. Early research on GNNs mainly focused on homophilous graphs, and recently there have been discussions whether specialized models are needed for the heterophilous setting [1, 32, 48, 24, 22, 33].
To measure the level of homophily, several homophily measures are used in the literature [1, 20, 32, 48], but these measures may significantly disagree with each other. In this work, we start by addressing the problem of how to properly measure the homophily level of a graph. For this, we formalize some desirable properties of a reasonable homophily measure and check which measures satisfy which properties. One essential property is called constant baseline and, informally speaking, it requires a measure to be not biased towards particular numbers of classes or their size balance. Our analysis reveals that commonly used homophily measures do not satisfy this property and thus cannot be compared across different datasets. In contrast, a measure that we call adjusted homophily (a.k.a. assortativity coefficient) satisfies most of the desirable properties while being rarely used in graph ML literature. Based on our theoretical analysis, we advise using adjusted homophily as a better alternative to the commonly used measures.
@@ 2.1 Popular homophily measures
Many GNN models implicitly make a so-called homophily assumption: that similar nodes are connected. Similarity can be considered in terms of node features or node labels. Usually, label homophily is analyzed, and we also focus on this aspect, leaving feature homophily for further studies. There are several commonly used homophily measures in the literature. Edge homophily [1, 48] is the fraction of edges that connect nodes of the same class:
@@
Node homophily [32] computes the fraction of neighbors that have the same class for all nodes and then averages these values across the nodes:
@@
These two measures are intuitive but have the downside of being sensitive to the number of classes and their balance, which makes them hard to interpret and incomparable across different datasets [20]. For example, suppose that each node in a graph is connected to one node of each class. Then, both edge homophily and node homophily for this graph will be equal to $\frac{1}{C}$. Thus, these metrics will produce widely different values for graphs with different numbers of classes, despite these graphs being similar in exhibiting no homophily. To fix these issues, Lim et al. [20] propose another homophily measure sometimes referred to as class homophily [22]. Class homophily measures excess homophily compared to a null model where edges are independent of the labels. More formally,
@@ 2.2 Desirable properties for homophily measures
Above, we discussed some disadvantages of existing homophily measures. In this section, we formalize and extend this discussion: we propose a list of properties desirable for a good homophily measure. Our analysis is motivated by recent studies of clustering and classification performance measures [8, 9], but not all their properties can be transferred to homophily measures. For instance, we do not require symmetry — a property that a measure does not change when we swap the compared objects — since homophily compares entities of different nature (a graph and a labeling). For the same reason, the distance property (requiring a measure to be linearly transformed to a metric distance) cannot be defined. On the other hand, some of our properties are novel.
@@
For instance, edge homophily and node homophily are empty class tolerant, while class homophily is not. Empty class tolerance is a new property that was not discussed in [8, 9] since classification and clustering evaluation measures are used within one given dataset, see Appendix B.3 for more details.
@@
In contrast to Gösgens et al. [8, 9], our notion of monotonicity requires the property to hold across graphs with different numbers of classes. This is caused by the empty class tolerance property.
@@ 2.4 Adjusted homophily
where we use the notation $\bar{p}(k)=\frac{D_{k}}{2|E|}$ (degree-weighted distribution of class labels). This measure is known in graph analysis literature as assortativity coefficient [28]. While assortativity is a general concept that is often applied to node degrees, it reduces to (2) when applied to categorical node attributes on undirected graphs. Unfortunately, this measure is rarely used in graph ML literature (Suresh et al. [41] is the only work we are aware of that uses it for measuring homophily), while our theoretical analysis shows that it satisfies many desirable properties. Indeed, the following theorem holds (see Appendix B.4 for the proof).
@@ 2.5 Edge-wise homophily vs classification evaluation measures
We conclude the analysis of homophily measures by establishing a connection between them and classification evaluation measures [8]. For this, let us first define edge-wise homophily measures. We say that a homophily measure is edge-wise if it is a function of the class adjacency matrix that we now define. Since we consider undirected graphs, each edge $\{u,v\}\in E$ gives two ordered pairs of nodes $(u,v)$ and $(v,u)$. We can define a class adjacency matrix $\mathcal{C}$ as follows: each matrix element $c_{ij}$ denotes the number of edges $(u,v)$ such that $y_{u}=i$ and $y_{v}=j$. Since the graph is undirected, the matrix $\mathcal{C}$ is symmetric. Note that monotonicity can be naturally put in terms of the class adjacency matrix: adding an edge between two nodes of the same class $i$ corresponds to incrementing the diagonal element $c_{ii}$ by two, and adding an edge between two nodes of different classes $i$ and $j$ corresponds to incrementing $c_{ij}$ and $c_{ji}$ by one.
@@
Interestingly, both Cohen’s Kappa and Matthews correlation coefficient correspond to adjusted homophily. As argued in [8], the Matthews coefficient is one of the best classification evaluation measures in terms of its theoretical properties. Our extended analysis confirms this conclusion for the corresponding homophily measure: we prove a stronger version of asymptotic constant baseline and also establish a stronger variant of monotonicity in the interval of large values of $h_{adj}$. The latter result is essential since in [8] it was only claimed that monotonicity is violated when $C>2$.
Another measure advised by [8] is symmetric balanced accuracy. Since in our case the class adjacency matrix is symmetric, it gives the same expression as balanced accuracy: $h_{bal}=\frac{1}{C}\sum_{k=1}^{C}\frac{|(u,v):y_{u}=y_{v}=k|}{D_{k}}.$ The obtained measure satisfies the maximal and minimal agreement properties. However, it is not empty class tolerant and thus is not monotone. The asymptotic constant baseline is also not satisfied: the value $c_{base}=1/C$ depends on the number of classes. Thus, despite this measure being suitable for classification evaluation, it cannot be used as a homophily measure. This difference is caused by the fact that homophily measures have to be comparable across datasets with different numbers of classes, while for classification evaluation it is not required.
@@
Finally, we also note that homophily measures can be directly related to community detection evaluation measures including the well-established characteristic in graph community detection literature called modularity [30]. See Appendix B.5 for a detailed discussion.
@@ 3 Label informativeness
To give an example, among strictly heterophilous graphs in which nodes never connect to nodes of the same class, there can be those where edges are drawn between particular pairs of classes (Figure 1(a)) and those where the class of a node cannot be derived from the class of its neighbor (Figure 1(b)). While adjusted homophily correctly captures the absence of homophily in these graphs, it is not designed to identify which type they belong to. However, distinguishing such graphs is practically important: informative neighbors can be very useful for models accounting for the graph structure.
We define a characteristic measuring the informativeness of a neighbor’s label for a node’s label. For example, in Figure 1(a), the neighbor’s label uniquely defines the node’s label. Thus, the node classification task is simple on this dataset, and we want our informativeness to be maximal for such graphs. Let us formalize this idea. Assume that we sample an edge $(\xi,\eta)\in E$ (from some distribution). The class labels of nodes $\xi$ and $\eta$ are then random variables $y_{\xi}$ and $y_{\eta}$. We want to measure the amount of knowledge the label $y_{\eta}$ gives for predicting $y_{\xi}$. The entropy $H(y_{\xi})$ measures the ‘hardness’ of predicting the label of $\xi$ without knowing $y_{\eta}$. Given $y_{\eta}$, this value is reduced to the conditional entropy $H(y_{\xi}|y_{\eta})$. In other words, $y_{\eta}$ reveals $I(y_{\xi},y_{\eta})=H(y_{\xi})-H(y_{\xi}|y_{\eta})$ information about the label. To make the obtained quantity comparable across different datasets, we say that label informativeness is the normalized mutual information of $y_{\xi}$ and $y_{\eta}$:
@@ 4.1 Characterizing real graph datasets
It is expected that datasets with high homophily $h_{adj}$ also have high $\mathrm{LI}$ since homophily implies informative neighbor classes. For medium-level homophily, $\mathrm{LI}$ can behave differently: for instance, ogbn-arxiv and twitter-hate have similar homophily levels, while the neighbors in ogbn-arxiv are significantly more informative. For heterophilous datasets, $\mathrm{LI}$ can potentially be very different, as we demonstrate in Section 4.2 on synthetic and semi-synthetic data. However, most existing heterophilous datasets have $\mathrm{LI}\approx 0$. This issue is partially addressed by new heterophilous datasets recently proposed in [33]. For the proposed roman-empire dataset, $\mathrm{LI}=0.11$ and $h_{adj}=-0.05$. Thus, while being heterophilous, this dataset has non-zero label informativeness, meaning that neighboring classes are somewhat informative. We believe that datasets with more interesting connectivity patterns will be collected in the future.
@@ 4.2 Correlation of LI with GNN performance
Recently, it has been shown that standard GNNs can sometimes perform well on non-homophilous datasets [22, 24]. We hypothesize that GNNs can learn more complex relationships between nodes than just homophily, and they will perform well as long as the node’s neighbors provide some information about this node. Thus, we expect $\mathrm{LI}$ to better correlate with the performance of GNNs than homophily. To illustrate this, we use carefully designed synthetic and semi-synthetic data. First, it allows us to cover all combinations of homophily levels and label informativeness. Second, we can control that only a connection pattern changes while other factors affecting the performance are fixed.
@@
To start with the most simple and controllable setting, we generate synthetic graphs via a variant of the stochastic block model (SBM) [13]. In this model, the nodes are divided into $C$ clusters, and for each pair of nodes $i,j$, we draw an edge between them with probability $p_{c(i),c(j)}$ independently of all other edges. Here $c(i)$ is a cluster assignment for a node $i$, which in our case corresponds to the node label $y_{i}$.
@@
We generated graphs according to the procedure described above with the expected node degree of 10 and various combinations of $p_{0},p_{1},p_{2}$. Given the class labels, the features are taken from the four largest classes in the cora dataset [38, 27, 46, 25]. We use the obtained graphs to train four popular GNN models: GCN [16], GraphSAGE [10], GAT [43], and Graph Transformer (GT) [40]. In total, we run more than 20000 experiments on synthetic datasets for each of the four models. A detailed description of data generation, training setup, and hyperparameters used is provided in Appendix F.1.
@@
Figure 2(a) shows the results for GraphSAGE. It can be seen that the performance is much more correlated with $\mathrm{LI}$ than with homophily. In particular, when $\mathrm{LI}$ is high, GraphSAGE achieves good performance even on strongly heterophilous graphs with negative homophily. We refer to Appendix F.2 for additional visualizations.
(a) Synthetic SBM graphs
(b) Semi-synthetic cora graphs from [24]
@@
Semi-synthetic data from [24]
Ma et al. [24] also argue that standard GNNs can perform well on certain heterophilous graphs. They construct semi-synthetic graphs by adding inter-class edges following different patterns to several real-world graphs, thus obtaining several sets of graphs with varying levels of homophily. Ma et al. [24] run experiments on these graphs and note that a standard GNN achieves strong performance on some heterophilous graphs.
Table 4: Spearman correlation between model accuracy and characteristics of semi-synthetic datasets from [24]
@@
We train GCN, GraphSAGE, GAT, and GT on the same modifications of the cora and citeseer graphs used in [24] and find that the models achieve strong performance when the graphs have high label informativeness. Training setup and hyperparameters used in these experiments are the same as above and are described in Appendix F.1. The performance of GraphSAGE on cora is shown in Figure 2(b). In Appendix F.3, we also show the performance of GraphSAGE on the citeseer dataset. Table 4 shows the Spearman correlation coefficients between accuracy and various homophily measures or $\mathrm{LI}$ for all models and both datasets. The correlation coefficients for homophily measures are all negative, while for $\mathrm{LI}$ they are positive and sufficiently large. This again confirms that $\mathrm{LI}$ indicates whether graph structure is helpful for GNNs.
Additionally, we conduct experiments with other datasets: Appendix F.4 describes the results for the LFR benchmark [17] and Appendix F.5 presents an experiment on synthetic data from [22].